What is the mathematical expression of this. A three-digit number whose hundreds digit is twice the tens digit and the tens digit is 3 more that the units digit
Answer by ikleyn(53751)

.

Let u be the units digit.

Then the tens digit is u+3,  

while the hundreds digit is 2*(u+3).



Notice that 2*(u+3) must be less or equal to 8, so

    2(u+3) <= 8,  so  u+3 <= 8/2 = 4,  which gives  0 <= u <= 1



Thus the formula for the value of the 3-digit number is


    100*2*(u+3) + 10*(u+3) + u = 200(u+3) + 10*(u+3) + u,


where  " u " is  EITHER  0  or  1.


If  u = 0,  then the 3-digit number is  630;

if  u = 1,  then the 3-digit number is  841.

Solved and explained.